Q1)If in a triangle, a= 6, b= 3 and cos(A-B)=4/5, then find its area?
Q2)Let the angles of a A,B,C of a triangle ABC be in A.P.and let b:c=under root(3):under root(2).Find the angle A?

Grade:11

1 Answers

Arjun Blouria
12 years ago

cos(A-B)=4/5

therefore,    1-tan2(A-B)/2  = 4/5 

                 1+tan2(A-B)/2

By componendo dividendo,

2tan2(A-B) =  5-4

     2             5+4